# LLMMapReduce-V3: Enabling Interactive In-Depth Survey Generation through a MCP-Driven Hierarchically Modular Agent System

### v1

@@ 1 Introduction
Generating high-quality long-form survey articles poses significant challenges to AI Agent systems Li et al. (2024); Zhang et al. (2025b); Qin et al. (2024). First, the system must effectively aggregate and synthesize information from a large collection of reference materials Nakano et al. (2021); Asai et al. (2023); Huang et al. (2024), which span various subdomains and perspectives. Second, it requires the ability to construct a coherent and comprehensive skeleton that guides the organization of content at a global level Wang et al. (2024); Wen et al. (2025). Third, as the upstream component, reference exploitation exerts a significant influence Elovic (2024), thus many AI-based survey generation tools have begun to focus on improving the quality of search results before the writing stage begins. Addressing these challenges demands controllable mechanisms for information acquisition, structural planning, and content generation.
To address these challenges, we propose LLMMapReduce-V3, an interactive and self-organized modular agents system for long-form survey generation. Our system builds upon the design of LLMMapReduce-V2 Wang et al. (2025), but extends it in several important ways to enable modularity, adaptability, and dynamic planning Qiu et al. (2025b, a). At the core of our approach is the use of the model context protocol (MCP), a standardized function-calling mechanism that allows tools and modules to be composed as independent MCP servers Raseed (2025); Unleash (2025). We proposes a multi-stage workflow of document digestion, skeleton construction and refinement, and survey writing. Building upon this foundation, we re-architect the system into a multi-agent paradigm, wherein core functionalities are decomposed across specialized agents, and the algorithmic workflow is encapsulated as a suite of tools in MCP servers Anthropic (2024) to facilitate agent-level invocation and coordination.
A key improvement in our system lies in pipeline flexibility with an agent planner Moura et al. (2024); Link-AGI (2024). Rather than following a fixed control flow, the agent receives a set of available MCP tools along with previous outputs, and dynamically selects the next module to invoke. This enables non-linear, adaptive workflows tailored to the specific needs of each writing task. To support user intent alignment, we also introduce human-in-the-loop interaction. Users begin by providing a target topic and basic writing instructions. The system then engages in a multi-turn dialogue to identify the user’s preferred and fine-grained research perspectives.
@@ 2.1 AI-Powered Automated Research
Early automated research systems focused on web-based information retrieval, with WebGPT Nakano et al. (2021) and Self-RAG Asai et al. (2023) pioneering LLM-based web browsing through human feedback or self-reflection mechanisms for adaptive retrieval and generation. Recent advances include sophisticated autonomous agents like GPT-Researcher Elovic (2024) with dual-agent architectures, ResearchAgent Huang et al. (2024) for iterative research idea generation through multi-agents collaboration, and products including Perplexity Deep Research AI (2024) and ChatGPT Deep Research OpenAI (2024). While some closed-source deep research systems can produce valuable results, we contend that flexible user involvement is critical for a truly practical system. We therefore propose an open-source, hierarchically modular system specifically designed to facilitate and support human intervention throughout the research process.
@@ 2.2 Survey Generation
LLM-driven survey generation has emerged with AutoSurvey Wang et al. (2024), which introduced a four-stage methodology that addresses context limitations and evaluation challenges, while InteractiveSurvey Wen et al. (2025) advances through personalized, interactive generation with continuous user customization of reference categorization and content synthesis. SurveyX Liang et al. (2025), on the other hand, focuses on extracting topics and content by pre-organizing the literature into the form of an attribute tree. Despite these advances, existing systems often lock users into a rigid, "all-or-nothing" paradigm. Stemming from inflexible integration approaches, they lack the necessary interfaces for iterative refinement and specialized customization. In response, we adopt MCP to modularize our algorithmic workflow, enabling the agents to iteratively select and invoke appropriate tools for survey generation.
@@ 2.3 Model Context Protocol and Modular Self-Organized Agent System
MCP Anthropic (2024) establishes open standards for connecting AI assistants to diverse tool sources through a unified client-server architecture. Recent applications demonstrate its potential for multi-agents intelligence Raseed (2025) and scalable systems Unleash (2025). Alita Qiu et al. (2025b) leverages MCP to autonomously construct and reuse external capabilities through task-related protocols, achieving scalable agentic reasoning with minimal pre-definition. AgentDistill Qiu et al. (2025a) enables training-free knowledge transfer via reusable MCP boxes, allowing smaller agents to achieve performance comparable to large LLM systems. These MCP-based self-evolution approaches demonstrate the its potential for adaptive agent systems. Our work first introduces the MCP-based hierarchically modular survey generation system, combining protocol standardization with specialized academic optimizations to enable unprecedented user customization and institutional integration while maintaining the quality of survey paper generation.
@@ 4.2 Agent Integration and Replaceable Invocation
| Manus AI Monica (2025) (Butterfly Effect AI) | ||||||
@@
| CoSearchAgent Chen et al. (2024) | ||||||
| OpenResearcher Liu et al. (2024) | ||||||
| Search-o1 Zhang et al. (2025a) | ||||||
| Agent-R1 Chen et al. (2025) | ||||||
@@
| CrewAI Moura et al. (2024) | ||||||
| AutoAgent Link-AGI (2024) | ||||||

### v2

@@ 1 Introduction
Generating high-quality long-form survey articles poses significant challenges to AI Agent systems Que et al. (2024); Wu et al. (2025); Wang et al. (2024b). First, the system must effectively aggregate and synthesize information from a large collection of reference materials Nakano et al. (2022); Asai et al. (2024); Baek et al. (2025), which span various subdomains and perspectives. Second, it requires the ability to construct a coherent and comprehensive skeleton that guides the organization of content at a global level Wang et al. (2024a); Wen et al. (2025). Third, as the upstream component, reference exploitation exerts a significant influence Elovic (2023), thus many AI-based survey generation tools have begun to focus on improving the quality of search results before the writing stage begins. Addressing these challenges demands controllable mechanisms for information acquisition, structural planning, and content generation.
To address these challenges, we propose LLMMapReduce-V3, an interactive and self-organized modular agents system for long-form survey generation. Our system builds upon the design of LLMMapReduce-V2 Wang et al. (2025), but extends it in several important ways to enable modularity, adaptability, and dynamic planning Qiu et al. (2025b, a). At the core of our approach is the use of the model context protocol (MCP), a standardized function-calling mechanism that allows tools and modules to be composed as independent MCP servers Basheer (2025); Pesenzon (2025). We proposes a multi-stage workflow of document digestion, skeleton construction and refinement, and survey writing. Building upon this foundation, we re-architect the system into a multi-agent paradigm, wherein core functionalities are decomposed across specialized agents, and the algorithmic workflow is encapsulated as a suite of tools in MCP servers Anthropic (2024) to facilitate agent-level invocation and coordination.
A key improvement in our system lies in pipeline flexibility with an agent planner CrewAIInc (2024); Chen et al. (2024). Rather than following a fixed control flow, the agent receives a set of available MCP tools along with previous outputs, and dynamically selects the next module to invoke. This enables non-linear, adaptive workflows tailored to the specific needs of each writing task. To support user intent alignment, we also introduce human-in-the-loop interaction. Users begin by providing a target topic and basic writing instructions. The system then engages in a multi-turn dialogue to identify the user’s preferred and fine-grained research perspectives.
@@ 2.1 AI-Powered Automated Research
Early automated research systems focused on web-based information retrieval, with WebGPT Nakano et al. (2022) and Self-RAG Asai et al. (2024) pioneering LLM-based web browsing through human feedback or self-reflection mechanisms for adaptive retrieval and generation. Recent advances include sophisticated autonomous agents like GPT-Researcher Elovic (2023) with dual-agent architectures, ResearchAgent Baek et al. (2025) for iterative research idea generation through multi-agents collaboration, and products including Perplexity Deep Research AI (2024) and ChatGPT Deep Research OpenAI (2024). While some closed-source deep research systems can produce valuable results, we contend that flexible user involvement is critical for a truly practical system. We therefore propose an open-source, hierarchically modular system specifically designed to facilitate and support human intervention throughout the research process.
@@ 2.2 Survey Generation
LLM-driven survey generation has emerged with AutoSurvey Wang et al. (2024a), which introduced a four-stage methodology that addresses context limitations and evaluation challenges, while InteractiveSurvey Wen et al. (2025) advances through personalized, interactive generation with continuous user customization of reference categorization and content synthesis. SurveyX Liang et al. (2025), on the other hand, focuses on extracting topics and content by pre-organizing the literature into the form of an attribute tree. Despite these advances, existing systems often lock users into a rigid, "all-or-nothing" paradigm. Stemming from inflexible integration approaches, they lack the necessary interfaces for iterative refinement and specialized customization. In response, we adopt MCP to modularize our algorithmic workflow, enabling the agents to iteratively select and invoke appropriate tools for survey generation.
@@ 2.3 Model Context Protocol and Modular Self-Organized Agent System
MCP Anthropic (2024) establishes open standards for connecting AI assistants to diverse tool sources through a unified client-server architecture. Recent applications demonstrate its potential for multi-agents intelligence Basheer (2025) and scalable systems Pesenzon (2025). Alita Qiu et al. (2025b) leverages MCP to autonomously construct and reuse external capabilities through task-related protocols, achieving scalable agentic reasoning with minimal pre-definition. AgentDistill Qiu et al. (2025a) enables training-free knowledge transfer via reusable MCP boxes, allowing smaller agents to achieve performance comparable to large LLM systems. These MCP-based self-evolution approaches demonstrate the its potential for adaptive agent systems. Our work first introduces the MCP-based hierarchically modular survey generation system, combining protocol standardization with specialized academic optimizations to enable unprecedented user customization and institutional integration while maintaining the quality of survey paper generation.
@@ 4.2 Agent Integration and Replaceable Invocation
| Manus AI Monica (Butterfly Effect AI) (2025) | ||||||
@@
| CoSearchAgent Gong et al. (2024) | ||||||
| OpenResearcher Zheng et al. (2024) | ||||||
| Search-o1 Li et al. (2025) | ||||||
| Agent-R1 Cheng et al. (2025) | ||||||
@@
| CrewAI CrewAIInc (2024) | ||||||
| AutoAgent Chen et al. (2024) | ||||||